# DisenSemi: Semi-supervised Graph Classification via Disentangled Representation Learning

### v1

@@ IV-B GNN-based Encoder
GNNs have recently emerged as effective approaches for learning graph-structured data. Prevailing GNNs for learning graph representations are based on the neural message passing among nodes to update their features and utilize a readout function to get the graph-level representation. However, we argue that these methods are holistic and neglect the hidden factors stemming from different aspects. In fact, multiple heterogeneous relations between nodes are mixed together and collapsed into one single edge, and these relations correspond to multiple aspects of discriminant features of the graph. Thus, given input graph $G_{i}$, we factorize it into $K$ factor graphs $\{G_{ik}\}_{k=1}^{K}$ to separately encode their features. Specifically, for each edge $e_{m}=\{x_{u},x_{v}\}$ in edge set $E(G_{i})=\{e_{1},\dots,e_{M}\}$, where $M$ is the total edge of graph $G_{i}$, we project the node feature $x_{v}\in\mathbb{R}^{d^{\prime}}$ with a transformation matrix $W\in\mathbb{R}^{d\times d^{\prime}}$ and generate the factor coefficients as follows:
@@
where $\Psi_{k}$ is the function that takes the features of node $u$ and $v$ as input to compute the attention score of edge $e_{m}$ under factor $k$, and the function can be implemented via a multi-layer perceptron (MLP). Following that, the attention score is normalized to get $E_{uvk}(G_{i})$, representing the coefficient of edge $e_{m}$ contained in the factor graph $G_{ik}$.
We represent each factor graph by its own edge coefficient $E_{uvk}(G_{i})$. By stacking $L$ message passing layers, the node embedding of each node $v$ in $G_{ik}$ is updated by recursively aggregating and combining its neighborhood features. Formally, the embedding of node $v$ in $l$-th layer is calculated as follows:
@@
where $N(v)$ denotes the neighborhood of node $v$, $\mathcal{A}_{k}^{(l)}$ represents the aggregation operation corresponding to the coefficient $E_{uvk}(G_{i})$ as edge weight for factor graph $G_{ik}$, and $\mathcal{C}_{k}^{(l)}$ represents the combination operations at layer $l$. In this paper, we adopt GraphConv [49] as our message passing layer due to its powerful expression ability and initialize the node embedding of each factor graph via a factor-specific transformation matrix $W_{K}\in\mathbb{R}^{\frac{d}{K}\times d^{\prime}}$.
Finally, for each factor graph $G_{ik}$, the graph-level representations can be attained by aggregating all node embeddings of the last $L$ layer with a readout function. Formally,
@@
where $V(G_{i})$ is the node set of $G_{i}$, READOUT can be a simple permutation invariant function such as the mean function. By considering all the hidden factors of graph data, we can get a factor-wise graph representation, namely, $Z_{G_{i}}=[z_{1}(G_{i}),\dots,z_{K}(G_{i})]$.
@@ IV-C Supervised Loss
With the GNN-based encoder in the supervised task, a factor-wise representation $Z_{SG_{i}}=[z_{S1}(G_{i}),\dots,z_{SK}(G_{i})]$ is extracted from labeled data, where different factor graphs will result in different features of the graph. Towards this end, we introduce $K$ learnable factor prototypes $\mathcal{C}=\{c_{k}\}_{k=1}^{K}$ to obtain the attention weight $\alpha_{G_{ik}}$ over latent factors for the input $G_{i}$:
@@
where $\phi$ is the similarity function, i.e., cosine similarity. After that, we weighted sum all latent factors to get the graph representation and fed it into a label prediction model to get the corresponding output, as shown in Fig. 3(b).
@@
where we adopt a two-layer MLP to map the graph representation extracted from different factor graphs to label predictions. For the supervised loss, we adopt cross-entropy, which can be defined as follows:
@@
where $y_{i}\in\mathbb{R}^{C}$ is the label with the one-hot encoding of the input graph data $G_{i}$, and $C$ corresponds to the number of graph categories in the dataset.
@@ IV-D MI-based Representation Disentanglement
For unsupervised learning, we can also get a factor-wise graph representation $Z_{UG_{i}}=[z_{U1}(G_{i}),\dots,z_{UK}(G_{i})]$ via the GNN-based encoder. Since some of the decomposed factor graphs may contain a similar structure without any other constrain, we propose several MI-based constraints to effectively disentangle the graph representation, as shown in Fig. 3(a).
@@ IV-D1 Intra-factor MI
where $I(z_{Uk}(G_{i}),h_{vk}^{(L)}(G_{i}))$ is the MI estimator modeled by discriminator $D_{k}$ and parameterized by a neural network. We use the binary cross-entropy (BCE) loss between samples from the joint (a.k.a., positive example) and product of marginals (a.k.a., negative examples) as the objective:
@@
where $h_{vk}^{(L)}(\tilde{G_{i}})$ is generated by the negative input graph $\tilde{G_{i}}$ sampled from a distribution $\tilde{p}=p$, which is identical to the empirical probability distribution of the input space. In practice, we could sample negative samples by an explicit (stochastic) corruption function or directly sample negative graph instances in a batch.
@@ IV-D2 Inter-factor MI
For graph representations $Z_{UG_{i}}$ extracted from different factor graphs, the MI between every two of them reaches its minimum value zero when $p(z_{Uk}(G_{i}),z_{Uk^{\prime}}(G_{i}))=p(z_{Uk}(G_{i}))p(z_{Uk^{\prime}}(G_{i}))$, which means that $z_{Uk}(G_{i})$ and $z_{Uk^{\prime}}(G_{i})$ are independent to each other. Thus, MI minimization among them encourages the factor-wise graph representation to learn different aspects of information for the graph. Recently, several MI upper bounds [50, 51] have been introduced for MI minimization. However, the estimation MI upper bound among these $K$ graph representations requires $K(K-1)$ times estimation, resulting in far more costs especially when $K$ is large. To alleviate this problem, as orthogonality can be seen as a special case of linear independence among representations, we loosen the constraint of minimization MI to orthogonality and the method has also been demonstrated to be effective by many previous studies [52]:
@@
Remark: In our model, we have also tested other MI minimization methods, such as Contrastive Log-ratio Upper Bound (CLUB) [51], which estimates MI by the difference of log-ratio of conditional distribution between positive and negative sample pairs. And these methods do not bring much change to the performance.
We maximize the MI between the factor graph and its corresponding graph representation to characterize each hidden factor while minimizing the MI among different factor graph representations for one input graph to enforce representation disentanglement. Formally, unsupervised loss of DisenSemi can be formulated as:
@@ IV-E Disentangled Consistency Regularization
For transferring the learned disentangled graph representations from the unsupervised encoder to the supervised encoder, we encourage representations extracted from two encoders with high MI when they correspond to the same factor of one graph data bridging supervised and unsupervised tasks. Specifically, we maximize NCE-based MI lower bound [47], which aims at making similar representations closer and different instances far away from each other. We treat each representation of a graph data instance as a distinct class and distinguish it from representations extracted from other graph data instances. Formally, given graph dataset $\mathcal{G}=\{G_{i}\}_{i=1}^{|\mathcal{G}^{L}|+|\mathcal{G}^{U}|}$, we assign each graph $G_{i}$ with a unique surrogate label $s_{i}=i$. Different from the ground-truth label $y_{i}$ of the labeled graph data, $s_{i}$ can be seen as the ID of the corresponding graph in the dataset $\mathcal{G}$. And in this way, the MI can be defined as:
@@
where $p(s_{i}|G_{i})$ corresponding to the identify discrimination over dataset, $p(k|G_{i})$ is the probability distribution of the $k$-th latent factors reflected in $G_{i}$, namely, can be $\alpha_{k}(G_{i})$. And $p(s_{i}|G_{i},k)$ corresponding to the identify discrimination subtask under $k$-th latent factor defined as:
@@
In effect, it is non-trivial to maximize the log-likelihood over the whole graph dataset due to the latent factors, where we need to calculate the posterior distribution defined with Bayes’ theorem as follows:
@@
Compared with prior distribution $p(k|G_{i})$ that inferred only given the graph $G_{i}$, $p(k|G_{i},s_{i})$ is the probability of the $k$-th latent factor effectively aligning both the supervised and unsupervised models when applied to the same graph data. However, it is prohibitive to compute posterior probability due to the term $p(s_{i}|G_{i},k)$ in Eq. 16, which needs all the instances in the dataset for computing the denominator. Therefore, we alternatively maximize the evidence lower bound (ELBO) of the log-likelihood, which is defined as:
@@
where $q(k|G_{i},s_{i})$ is a variational distribution to approximate the posterior probability $p(k|G_{i},s_{i})$. Here, we formalize the variational distribution by:
@@
Notice that the process is a variation of the Variational EM algorithm, where we infer $q(k|G_{i},s_{i})$ at the E-step and optimize the ELBO at the M-step.
@@ IV-F2 Complexity Analysis
To ensure scalability with large-scale datasets, we leverage mini-batch with size $B=B^{L}+B^{U}$ to compute gradients. The computational consumption is mainly composed of three parts: (i) the unsupervised module; (ii) the supervised module; (iii) the disentangled consistency regularization. Given the graph with an average number of nodes $|V|$ and edges $|E|$, the number of GNN layer and factor graphs is $L$ and $K$, and the representation dimension is $d$. For (i) and (ii), the time complexity of the GNN-based encoder $O(|E|Ld)$. For (i), the additional computational complexity of intra-factor and inter-factor MI constraints is $O(|V|d)$ and $O((K-1)d/2)$ respectively. For (ii), the additional computational complexity of the label prediction model is $O(d+(d/K)^{2}+Cd/K)$ with $C$ classes in the task. For (iii), we perform disentangled consistency regularization within minibatch, which takes $O(Bd)$ for each graph. To sum up, we have the overall complexity of DisenSemi, $O((2|E|L+|V|+(K+1)/2+d/K^{2}+C/K+B)d)$, which scales linearly in terms of the number of nodes and edges in each graph instance. Moreover, DisenSemi requires $K$ factor graphs, prototypes and coreresponding GNN-based encoder in both supervised and unsupervised module, the space complexity is $O((|V|+|E|)Ld+d+K|E|)$.
@@ V Experiment
We introduce the experimental settings and conduct extensive experiments on ten real-world benchmark datasets to validate the effectiveness of our proposed DisenSemi. We focus on answering the following research questions:
  • RQ1: How does our proposed model DisenSemi perform compared with other state-of-the-art models for semi-supervised graph classification?
  • RQ2: Is it necessary to keep key components of DisenSemi (i.e. GNN-based encoder, MI-based representation disentanglement and consistency regularization)?
  • RQ3: How do different hyper-parameters in DisenSemi impact the performance?
@@ V-A1 Datasets and Baselines
following the previous works [16, 14], which includes four molecule datasets: MUTAG, PTC-MR, NCI1 and OGB-HIV, one bioinformatic dataset: PROTEINS and five social network datasets: IMDB-BINARY (IMDB-B), IMDB-MULTI (IMDB-M), REDDIT-BINARY (REDDIT-B), REDDIT-MULTI-5K (REDDIT-M5K) and COLLAB. Following recent works [16, 15], we adopt one-hot degree and centrality features instead when node attributes are not available in the datasets. We compare our proposed DisenSemi with sixteen recent proposed baselines including seven traditional graph classification approaches: GK [10], SP [7], WL [8], MLG [9], DGK [54], Sub2Vec [55] and Graph2Vec [56] and nine GNN-based semi-supervised graph classification methods: MVGRL [57], GraphCL [15], JOAO [17], DGCL [27], RGCL [58], InfoGraph [16], GLA [14], TGNN [18], GraphSpa [20].
@@ V-A3 Implementation Details
We implement our DisenSemi model in Pytorch. For all datasets, the embedding size of DisenSemi and other baselines (except GK, SP, WL, MLG) is fixed to 128. Since the ground-truth number of the latent factors is unknown, we search the number of factor graphs $K$ amongst $\{2^{0},2^{1},2^{2},2^{3},2^{4}\}$, corresponding embedding dimensions are $\{128,64,32,16,8\}$ per factor graph. We set three message passing layers for our method and the hyper-parameters $\lambda=0.001,\gamma=0.001$. We optimize DisenSemi with Adam optimizer by setting the learning rate to 0.005 and the number of epochs is set to 200 for all the methods. For all the traditional graph classification approaches, we compute similarities between graphs and feed them into a downstream Support Vector Machines (SVM) classifier to evaluate the classification accuracy. For MVGRL, GraphCL, JOVO, DGCL and RGCL, we pre-train the model with unlabeled graphs first and fine-tune the model with the labeled graphs. For InfoGraph, GLA, TGNN, GraphSpa and our proposed DisenSemi, we integrate the pre-training and fine-tuning phases together. The source code of DisenSemi is available at https://github.com/jamesyifan/DisenSemi.

### v2

@@ IV-B GNN-based Encoder
GNNs have recently become powerful tools for learning from graph-structured data. Current GNNs typically use neural message passing to update node features and apply a readout function to obtain graph-level representations. However, we argue that these methods are holistic and neglect the hidden factors stemming from different aspects. In fact, multiple relations between nodes are always mixed together and represented as one single edge, and these relations correspond to distinct discriminant aspects of the graph. Thus, given input graph $G_{i}$, we factorize it into $K$ factor graphs $\{G_{ik}\}_{k=1}^{K}$ to separately encode their features. Specifically, for each edge $e_{m}=\{x_{u},x_{v}\}$ in edge set $E(G_{i})=\{e_{1},\dots,e_{M}\}$, where $M$ is the total edge of graph $G_{i}$, we project the node feature $x_{v}\in\mathbb{R}^{d^{\prime}}$ with a transformation matrix $W\in\mathbb{R}^{d\times d^{\prime}}$ and calculate the factor coefficients as:
@@
where $\Psi_{k}$ is the function that computes the attention score for edge $e_{m}$ under factor $k$ with the features of nodes $u$ and $v$ as input, the function can be implemented via a multi-layer perceptron (MLP). Following that, the attention score is normalized to get $E_{uvk}(G_{i})$, representing the coefficient of edge $e_{m}$ contained in the factor graph $G_{ik}$.
We represent each factor graph by its own edge coefficient $E_{uvk}(G_{i})$. By stacking $L$ message passing layers, the node embedding of each node $v$ in $G_{ik}$ is updated by recursively aggregating and combining its neighborhood features. Formally, the embedding of node $v$ at the $l$-th layer can be:
@@
where $N(v)$ represents the neighborhood of node $v$, $\mathcal{A}_{k}^{(l)}$ represents the aggregation operation corresponding to the coefficient $E_{uvk}(G_{i})$ as edge weight for factor graph $G_{ik}$, and $\mathcal{C}_{k}^{(l)}$ represents the combination operations at layer $l$. In this paper, we adopt GraphConv [44] as our message passing layer due to its strong expressive power and initialize the node embedding of each factor graph via a factor-specific transformation matrix $W_{K}\in\mathbb{R}^{\frac{d}{K}\times d^{\prime}}$.
Finally, for each factor graph $G_{ik}$, the graph-level representations are derived by aggregating the embeddings of all nodes from the last $L$-th layer through a readout function. Formally,
@@
where $V(G_{i})$ denotes the node set of $G_{i}$, READOUT can be implemented as a simple permutation invariant function (i.e., mean function). By considering all the hidden factors of graph data, we can get a factor-wise graph representation, namely, $Z_{G_{i}}=[z_{1}(G_{i}),\dots,z_{K}(G_{i})]$.
@@ IV-C Supervised Loss
With the GNN-based encoder in the supervised task, a factor-wise representation $Z_{SG_{i}}=[z_{S1}(G_{i}),\dots,z_{SK}(G_{i})]$ is extracted from labeled data, where different factor graphs will result in different features of the graph. Towards this end, we define $K$ learnable prototypes $\mathcal{C}=\{c_{k}\}_{k=1}^{K}$ to obtain the attention weight $\alpha_{G_{ik}}$ of factor $k$ for the input $G_{i}$:
@@
where $\phi$ denotes the similarity function (i.e., cosine similarity). After that, we weighted sum all latent factors to get the graph representation and fed it into a prediction model to get the output label. The process is shown in Fig. 2(b).
@@
where we adopt a two-layer MLP to map the graph representation extracted from different factor graphs to label predictions. For the supervised loss, we adopt cross-entropy, defined as:
@@
where $y_{i}\in\mathbb{R}^{C}$ is the label with the one-hot encoding for the graph $G_{i}$, $C$ corresponds to the total number of classes in the graph dataset.
@@ IV-D MI-based Representation Disentanglement
For unsupervised learning, we can also get a factor-wise graph representation $Z_{UG_{i}}=[z_{U1}(G_{i}),\dots,z_{UK}(G_{i})]$ via the GNN-based encoder. Since some of the decomposed factor graphs may exhibit similar structures without additional constraints, we propose several MI-based constraints to effectively disentangle the graph representation, as shown in Fig. 2(a).
@@ IV-D1 Intra-factor MI
where $I(z_{Uk}(G_{i}),h_{vk}^{(L)}(G_{i}))$ is the MI estimator, modeled by a discriminator $D_{k}$ and implemented through a neural network. We use the binary cross-entropy (BCE) loss between positive samples from the joint distribution and negative samples from the product of the marginals as the objective:
@@
where $h_{vk}^{(L)}(\tilde{G_{i}})$ is generated by the negative input graph $\tilde{G_{i}}$. We sample $\tilde{G_{i}}$ from a distribution $\tilde{p}=p$ to match the empirical distribution of the input space. In practice, we could sample negative samples by an explicit (stochastic) corruption function or directly sample negative graph instances in a batch.
@@ IV-D2 Inter-factor MI
For graph representations $Z_{UG_{i}}$ extracted from different factor graphs, the MI between every two of them reaches its minimum value zero when $p(z_{Uk}(G_{i}),z_{Uk^{\prime}}(G_{i}))=p(z_{Uk}(G_{i}))p(z_{Uk^{\prime}}(G_{i}))$, which means that $z_{Uk}(G_{i})$ and $z_{Uk^{\prime}}(G_{i})$ are independent to each other. Thus, MI minimization among them encourages the factor-wise graph representation to learn different aspects of information for the graph. Recently, several MI upper bounds [45, 46] have been introduced for MI minimization. However, the estimation MI upper bound among these $K$ graph representations requires $K(K-1)$ times estimation, resulting in far more costs especially when $K$ is large. To alleviate this problem, as orthogonality is a specific instance of linear independence, we loosen the constraint of minimization MI to orthogonality and the method has also been demonstrated to be effective by many previous studies [47]:
@@
Remark: We have also tested other MI minimization methods, such as Contrastive Log-ratio Upper Bound (CLUB) [46], which estimates MI by the log-ratio difference between the conditional distribution of positive and negative pairs. And these methods do not significantly impact the performance.
We maximize the MI between the factor graph and its corresponding graph representation to characterize each hidden factor while minimizing the MI among different factor graph representations to enforce representation disentanglement. Formally, unsupervised loss of DisenSemi can be:
@@ IV-E Disentangled Consistency Regularization
For transferring the learned disentangled graph representations from the unsupervised encoder to the supervised encoder, we encourage representations extracted from two encoders with high MI when they correspond to the same factor of one graph data bridging supervised and unsupervised models. Specifically, we maximize NCE-based MI lower bound [42], which seeks to bring similar representations closer together while pushing dissimilar instances further apart. We treat each representation of a graph data instance as a distinct class and distinguish it from representations extracted from other graph data instances. Formally, given graph dataset $\mathcal{G}=\{G_{i}\}_{i=1}^{|\mathcal{G}^{L}|+|\mathcal{G}^{U}|}$, we assign each graph $G_{i}$ with a unique surrogate label $s_{i}=i$. Unlike the label $y_{i}$ contained in labeled graph, $s_{i}$ can be seen as the corresponding ID of the instance in the dataset $\mathcal{G}$. In this way, the MI can be defined as:
@@
where $p(s_{i}|G_{i})$ corresponding to the identify discrimination over dataset, $p(k|G_{i})$ denotes the probability distribution that $G_{i}$ is reflected via the $k$-th latent factor, namely, can be $\alpha_{k}(G_{i})$. And $p(s_{i}|G_{i},k)$ denotes to the discrimination subtask to identify the corresponding ID under the $k$-th latent factor, which can be defined as:
@@
In practice, maximizing the log-likelihood across the entire graph dataset is difficult due to the latent factors. Therefore, we calculate the posterior distribution with Bayes’ theorem, which can be defined as:
@@
Compared with prior distribution $p(k|G_{i})$ that inferred only given the graph $G_{i}$, $p(k|G_{i},s_{i})$ is the probability that reflects how well the $k$-th latent factor aligns both supervised and unsupervised models when applied to the same graph data. However, computing the posterior probability is prohibitive due to the term $p(s_{i}|G_{i},k)$ in Eq. 16, which requires considering all instances in the dataset to compute the denominator. Therefore, we alternatively maximize the evidence lower bound (ELBO) of the log-likelihood, which is defined as:
@@
where $q(k|G_{i},s_{i})$ denotes a variational distribution that approximates the posterior probability $p(k|G_{i},s_{i})$. Here, we formalize the variational distribution by:
@@
Notice that the process is a variant of the Variational EM algorithm, where $q(k|G_{i},s_{i})$ is inferred during the E-step, and the ELBO is optimized during the M-step.
@@ IV-F2 Complexity Analysis
To ensure scalability with large-scale datasets, we leverage mini-batch with size $B=B^{L}+B^{U}$ to compute gradients. The computational consumption is mainly composed of three parts: (i) the unsupervised module; (ii) the supervised module; (iii) the disentangled consistency regularization. Given the graph with an average number of nodes $|V|$ and edges $|E|$, the number of GNN layer and factor graphs is $L$ and $K$, and the representation dimension is $d$. For (i) and (ii), the time complexity of the GNN-based encoder $O(|E|Ld)$. For (i), the additional computational complexity of intra-factor and inter-factor MI constraints is $O(|V|d)$ and $O((K-1)d/2)$ respectively. For (ii), the additional computational complexity of the label prediction model is $O(d+(d/K)^{2}+Cd/K)$ with $C$ classes in the task. For (iii), we perform disentangled consistency regularization within minibatch, which takes $O(Bd)$ for each graph. To summarize, we have the overall complexity of DisenSemi, $O((2|E|L+|V|+(K+1)/2+d/K^{2}+C/K+B)d)$, which scales linearly w.r.t. the number of nodes and edges in each graph. Moreover, DisenSemi requires $K$ factor graphs, prototypes and coreresponding GNN-based encoder in both supervised and unsupervised module, the space complexity is $O((|V|+|E|)Ld+d+K|E|)$.
@@ V Experiment
We describe the experimental setup and present a series of extensive experiments conducted on ten real-world benchmark datasets to evaluate the performance of our proposed method. Our goal is to address the following research questions:
  • RQ1: How does our proposed method compare to other advanced models in semi-supervised graph classification?
  • RQ2: Are the key components of our method (GNN-based encoder, MI-based representation disentanglement, and consistency regularization) essential?
  • RQ3: How do different hyper-parameters affect the performance of our method?
@@ V-A1 Datasets and Baselines
following the previous works [16, 14], which includes four molecule datasets: MUTAG, PTC-MR, NCI1 and OGB-HIV, one bioinformatic dataset: PROTEINS and five social network datasets: IMDB-BINARY (IMDB-B), IMDB-MULTI (IMDB-M), REDDIT-BINARY (REDDIT-B), REDDIT-MULTI-5K (REDDIT-M5K) and COLLAB. Following recent works [16, 15], we adopt one-hot degree and centrality features instead when the datasets do not include node attributes. We compare our proposed DisenSemi with sixteen recent proposed baselines including seven traditional graph classification approaches: GK [10], SP [7], WL [8], MLG [9], DGK [49], Sub2Vec [50] and Graph2Vec [51] and nine GNN-based semi-supervised graph classification methods: MVGRL [52], GraphCL [15], JOAO [17], DGCL [26], RGCL [53], InfoGraph [16], GLA [14], TGNN [18], GraphSpa [19].
@@ V-A3 Implementation Details
For all datasets, the embedding size of DisenSemi and other baselines (except GK, SP, WL, MLG) is fixed to 128. As the exact number of latent factors is unknown, we test different values for the number of factor graphs $K$ from the set $\{2^{0},2^{1},2^{2},2^{3},2^{4}\}$, corresponding embedding dimensions are $\{128,64,32,16,8\}$ per factor graph. We set three message passing layers for our method and the hyper-parameters $\lambda=0.001,\gamma=0.001$. The training is conducted for 200 epochs across all methods. For all the traditional graph classification approaches, we compute similarities between graphs and feed them into a downstream Support Vector Machines (SVM) classifier to evaluate the classification accuracy. For MVGRL, GraphCL, JOVO, DGCL and RGCL, we first pre-train the model with unlabeled graphs first and then fine-tune it with labeled graphs. For InfoGraph, GLA, TGNN, GraphSpa and our proposed DisenSemi, we integrate the pre-training and fine-tuning phases together. The source code of DisenSemi can be found at https://github.com/jamesyifan/DisenSemi.